# TS-RAG: Retrieval-Augmented Generation based Time Series Foundation Models are Stronger Zero-Shot Forecaster

### v3

@@ A.4 Details of Inference Datasets
We experiment the zero-shot forecasting on the widely adopted Electricity Transformer Temperature (ETT) datasets [10], Weather, Electricity [50], and Exchange Rate from  [51]. ETT datasets are comprised of roughly two years of data from two locations in China. The data are further divided into four distinct datasets, each with different sampling rates: ETTh1 and ETTh2 are sampled hourly, and ETTm1 and ETTm2 are sampled every 15 minutes. Every ETT dataset includes six power load features and a target variable: the oil temperature. The Electricity dataset comprises records of electricity consumption from 321 customers and is measured with a 1-hour sampling rate. The Weather dataset contains one-year records from 21 meteorological stations located in Germany. The sampling rate for the Weather dataset is 10 minutes. The Exchange Rate dataset includes the daily exchange rates of eight foreign countries, including Australia, Britain, Canada, Switzerland, China, Japan, New Zealand, and Singapore, ranging from 1990 to 2016.
@@ A.5 Evaluation Metrics
[…]
@@ A.6 Efficiency Analysis
At inference time, TS-RAG introduces a top-$k$ retrieval step. We use FAISS to perform fast nearest-neighbor search over the retrieval knowledge base. On ETTh, retrieval adds 9.2 ms latency per query; the ARM-augmented forward pass adds 0.44 ms. Total inference time is 9.62 ms/query. This overhead is minor and offset by consistent gains (e.g., --6.84% MSE on Exchange rate), making TS-RAG suitable for real-time use. Notably, compared to RAF, which requires 3.47 seconds per query, TS-RAG achieves over 360$\times$×faster inference speed, highlighting its superior efficiency and practicality.
@@ B.1 Generalization across Backbones
To verify the generalization ability of TS-RAG, we further evaluate it on the MOMENT [37] backbone, in addition to the Chronos-bolt results reported in the main text. The same retrieval-augmented framework is applied without modifying the backbone architecture or training recipes, and all retrieval settings follow those used with Chronos-bolt for fair comparison. Note that to enable the MOMENT model to perform zero-shot long-term forecasting, we pretrain a prediction head on the same pretraining data as TS-RAG. Table 4 reports the results on both backbones. TS-RAG consistently improves zero-shot forecasting performance regardless of the backbone, demonstrating its plug-and-play nature.
@@ B.4 Effect of Retrieval Lookback Length
B.5 Impact of Retriever Encoder Choice
@@
B.6 Comparion between TS-RAG and RAF
@@ C.1 Case Studies on Retrieval Effectiveness
Figures 6 and  7 illustrate the retrieval performance of TS-RAG on the ETTh1 and ETTh2 datasets. The retrieval results demonstrate that TS-RAG effectively identifies historical patterns with strong structural similarity to the input, particularly in terms of periodicity and trend dynamics. In ETTh1, the retrieved sequences capture complex fluctuations and local variations, aligning well with the seasonal patterns of the input. Meanwhile, in ETTh2, where the time series exhibits smoother periodicity, the retrieved sequences show almost perfect alignment, indicating the presence of highly consistent cyclic behavior. These results suggest that retrieval augmentation enhances forecasting by leveraging time series patterns that closely match the current context, particularly in datasets with strong seasonal dependencies.
@@ C.2 Case Studies on Retrieval-Augmented Forecasting
Figures 8 and  9 showcase the impact of retrieval augmentation on forecasting accuracy in the Weather dataset. Figure 8 highlights a situation where the baseline TSFM struggles to capture a sudden trend shift, leading to a significant forecasting error. By incorporating retrieved forecasting horizons, TS-RAG successfully adapts to the trend change. Figure 9 demonstrates how retrieval augmentation enhances peak prediction. The standard TSFM underestimates the upcoming peak, whereas TS-RAG, guided by similar retrieved patterns, generates a more accurate forecast. These case studies illustrate how retrieval-augmented forecasting helps models better adapt to complex temporal patterns, improving robustness in real-world forecasting tasks.
@@
Appendix D Limitations
@@
Appendix E Boarder Impact

### v4

@@ A.4 Details of Inference Datasets
We experiment the zero-shot forecasting on the widely adopted Electricity Transformer Temperature (ETT) datasets (zhou2021informer), Weather, Electricity (wu2023timesnet), and Exchange Rate from  (lai2018modeling). ETT datasets are comprised of roughly two years of data from two locations in China. The data are further divided into four distinct datasets, each with different sampling rates: ETTh1 and ETTh2 are sampled hourly, and ETTm1 and ETTm2 are sampled every 15 minutes. Every ETT dataset includes six power load features and a target variable: the oil temperature. The Electricity dataset comprises records of electricity consumption from 321 customers and is measured with a 1-hour sampling rate. The Weather dataset contains one-year records from 21 meteorological stations located in Germany. The sampling rate for the Weather dataset is 10 minutes. The Exchange Rate dataset includes the daily exchange rates of eight foreign countries, including Australia, Britain, Canada, Switzerland, China, Japan, New Zealand, and Singapore, ranging from 1990 to 2016.
@@ A.5 Evaluation Metrics
$\text{ MSE }=\frac{1}{H}\sum_{h=1}^{T}\left(\mathbf{Y}_{h}-\hat{\mathbf{Y}}_{h}\right)^{2},\qquad\text{ MAE }=\frac{1}{H}\sum_{h=1}^{H}\left|\mathbf{Y}_{h}-\hat{\mathbf{Y}}_{h}\right|,$
@@ A.6 Efficiency Analysis
[…]
@@ B.1 Generalization across Backbones
To verify the generalization ability of TS-RAG, we further evaluate it on the MOMENT (goswami2024moment) backbone, in addition to the Chronos-Bolt results reported in the main text. The same retrieval-augmented framework is applied without modifying the backbone architecture or training recipes, and all retrieval settings follow those used with Chronos-Bolt for fair comparison. Note that to enable the MOMENT model to perform zero-shot long-term forecasting, we pretrain a prediction head on the same pretraining data as TS-RAG. Table 5 reports the results on both backbones. TS-RAG consistently improves zero-shot forecasting performance regardless of the backbone, demonstrating its plug-and-play nature.
@@
Backbone Metric ETTh1 ETTh2 ETTm1 ETTm2 Weather Electricity Exchange Average
@@ B.3 Sensitivity to the Number of Retrieved Sequences
B.4 Impact of Retriever Encoder Choice
@@
Table 11: Zero-shot forecasting results under different retrieval distance metrics. Both MSE and MAE are reported.
Dataset w/o RAG Euclidean Cosine DTW
MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 0.3616 0.3650 0.3557 0.3624 0.3558 0.3624 0.3606 0.3647
ETTh2 0.2517 0.2992 0.2451 0.2982 0.2465 0.2982 0.2511 0.3001
ETTm1 0.3109 0.3185 0.2906 0.3114 0.2906 0.3111 0.3100 0.3183
ETTm2 0.1487 0.2236 0.1466 0.2231 0.1458 0.2224 0.1489 0.2240
Exchange 0.0673 0.1780 0.0627 0.1718 0.0624 0.1716 0.0665 0.1776
B.5 Effect of Retrieval Distance Metrics
To further investigate the effect of retrieval metrics, we conduct additional ablation studies using cosine similarity and DTW distance. For a fair comparison, cosine similarity is calculated over the embeddings generated by the Chronos encoder, while the DTW distance is computed directly in the original time-series space.
From the results in Table 11, we observe that Euclidean and cosine distances achieve very similar performance, both clearly outperforming the baseline setting (i.e., without RAG). In contrast, the DTW distance only achieves modest improvement over the baseline, and in some cases even performs slightly worse. This indicates that embedding-based retrieval methods are more effective and robust for identifying relevant forecasting references than distance measures computed directly in the raw time-series space. Notably, this finding also aligns with prior observations liu2024retrievaldiffusion that correlation-based methods are significantly inferior to embedding-based methods for retrieving forecasting references.
B.6 Comparison between TS-RAG and RAF
@@ C.1 Case Studies on Retrieval Effectiveness
Figures 7 and  8 illustrate the retrieval performance of TS-RAG on the ETTh1 and ETTh2 datasets. The retrieval results demonstrate that TS-RAG effectively identifies historical patterns with strong structural similarity to the input, particularly in terms of periodicity and trend dynamics. In ETTh1, the retrieved sequences capture complex fluctuations and local variations, aligning well with the seasonal patterns of the input. Meanwhile, in ETTh2, where the time series exhibits smoother periodicity, the retrieved sequences show almost perfect alignment, indicating the presence of highly consistent cyclic behavior. These results suggest that retrieval augmentation enhances forecasting by leveraging time series patterns that closely match the current context, particularly in datasets with strong seasonal dependencies.
@@ C.2 Case Studies on Retrieval-Augmented Forecasting
Figures 9 and  10 showcase the impact of retrieval augmentation on forecasting accuracy in the Weather dataset. Figure 9 highlights a situation where the baseline TSFM struggles to capture a sudden trend shift, leading to a significant forecasting error. By incorporating retrieved forecasting horizons, TS-RAG successfully adapts to the trend change. Figure 10 demonstrates how retrieval augmentation enhances peak prediction. The standard TSFM underestimates the upcoming peak, whereas TS-RAG, guided by similar retrieved patterns, generates a more accurate forecast. These case studies illustrate how retrieval-augmented forecasting helps models better adapt to complex temporal patterns, improving robustness in real-world forecasting tasks.
@@
Appendix D Analysis on Time Series Characteristics Influencing TS-RAG Effectiveness
We further investigate which characteristics of time series data make TS-RAG more effective. Specifically, we analyze four statistical properties of the benchmark datasets: autocorrelation, noise ratio, volatility, and stationarity, and study how they correlate with TS-RAG’s performance improvement over its baseline TSFM.
  • Autocorrelation measures the strength of temporal dependencies, i.e., how strongly current observations relate to their past values.
  • Noise ratio reflects the level of irregular fluctuations in the series, computed as the variance ratio of first differences.
  • Volatility quantifies the variability of the series relative to its mean, computed as the standard deviation divided by the mean value of each sequence. This ratio reflects local fluctuation intensity.
  • Stationarity estimates how stable the distribution of a time series is over time, approximated by the variance of first differences.
We compute these characteristics for all samples in each dataset and report their average values to represent dataset-level properties. We then calculate the Pearson correlation between each characteristic and the MSE improvement of TS-RAG over the baseline model. Results are summarized in Table 13.
Table 13: Quantitative analysis of dataset characteristics and their correlation with TS-RAG performance gains.
Dataset Autocorr. Noise Ratio Volatility Stationarity MSE Diff
ETTh1 0.7799 0.4391 3.1655 0.1752 0.0059
ETTh2 0.6070 0.7217 -0.5386 0.0843 0.0066
ETTm1 0.8437 0.2915 -0.9014 0.0536 0.0203
ETTm2 0.6848 0.4480 17.9827 0.0348 0.0021
Correlation 0.70 -0.55 -0.65 -0.19
Our analysis shows that datasets with stronger autocorrelation tend to benefit more from TS-RAG, whereas higher noise levels and greater volatility correspond to smaller improvements. This suggests that the retrieval-augmentation mechanism is particularly effective when temporal dependencies are strong and the underlying patterns are relatively stable.
Appendix E Limitations
@@
Appendix F Broader Impact